1. Price Elasticity of Demand. Before: P=1, Qd = 4; After: P =3; Qd =0.

Given your answer to #1, is this demand curve elastic or inelastic over this range? Why?
Answer #1

P1 = 1 QD1 = 4

P2 = 3 QD2= 0

Price Elasticity of demand = %age change in Quantity demanded / %age change in Price

%age change in quantity demanded = QD2 - QD 1 / QD1 * 100

= 0 - 4/ 4 * 100 = -100%

%age change in Price = P2 - P1 / P1 * 100

= 3 - 1 / 1 * 100 = 200%

Now PED = 100% / 200% = 0.5

As the proportionate change in Qunatity demanded is less than Proportionate change in Price the demand curve is said to be inelastic as Price elasticity of demand for good is below 1